[Koji and Kiwamu]
We did some more vacuum works today. It is getting ready to pump down.
(what we did)
- alignment of the POY mirrors. Now the beam is coming out from the ITMY chamber successfully
- leveling of the tables (except for the IOO and OMC chamber)
- realigned the beam axis down to the Y arm because the leveling of the BS table changed the alignments.
- installed IP_POS mirrors
- aligned the green beam and made it overlap with IR beam path.
- repositioned green steering mirrors since one of them are too close to the dark beam path